# A Near-optimal, Scalable and Corruption-tolerant Framework for Stochastic Bandits: From Single-Agent to Multi-Agent and Beyond

### v1

@@ B.1 Lemmas for Proving Theorem 2
Lemma 8
For any epoch $m$, all arms $k\neq k_{m}$ must satisfy $\widetilde{n}_{k}^{m}\leq n_{k}^{m}$ and $\hat{n}_{k}^{m}\geq n_{k}^{m}$.
Proof 8
@@
For each arm $k$ which hasn’t a self-loop, since $\widetilde{n}_{k_{m}}^{m}\geq 2^{-m}\geq n_{k}^{m}$ for all arms $k\neq k_{m}$, so we complete the proof.
Same as the proof in Appendix A, we only need to change $\widetilde{n}_{k}^{m}$ to $\hat{n}_{k}^{m}$ to obtain the following lemmas.
Lemma 9
@@
We also can always guarantee the following inequality:
@@
$-\frac{3}{7}\Delta-\frac{3}{4}2^{-m}-6\rho_{m}\leq\Delta_{k^{*}}^{m}\leq\frac{1}{7}\Delta+2^{-(m-1)}+2\rho_{m}.$
Proof 9
@@
$-\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\hat{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\hat{n}_{k}^{m}}}.$
@@
$-\frac{D_{m}}{N_{m}}-\frac{1}{4}\max\{\Delta_{k^{*}}^{m-1},\Delta^{m-1}\}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{N_{m}}.$
@@
Assuming the statement is true for $m-1$, for arm $k\neq k^{*}$, since there are at least two arms: one optimal arm and one sub-optimal arm, we can obtain
@@
This proof is complete.
Lemma 12
For any strongly observable directed graph $G$ with the independence number $\alpha$, the obtained out-dominating set ${\mathcal{D}}$ must satisfies $|{\mathcal{D}}|\leq\alpha(1+2\ln(K/\alpha))$ by Algorithm 3. Especially, when $G$ is a acyclic graph (include undirected graphs), we have $|{\mathcal{D}}|\leq\alpha$.
Proof 10
Recalling the definition of the no-root vertex, for any strongly observable directed graph $G_{1}$ with the independence number $\alpha$, we can get that if we remove an no-root vertex $k$ and its out-degree neighbors, the remaining graph $G_{2}$ whose independence number at most $\alpha-1$. That’s because no vertex in graph $G_{2}$ is connected to the vertex $k$. By the definition of strongly observable directed acyclic graph, we can always get the no-root vertex for the remaining graph, which means that $|{\mathcal{D}}|\leq\alpha$.
By Lemma 43 in Dann et al. [2023], for any strongly observable directed graph $G_{s}=(V_{s},E)$ with the independence number $\alpha$, we set $p_{1}=p_{2}=\cdots=p_{|V_{s}|}=\frac{1}{|V_{s}|}$, which can derive the following inequality:
$\exists\quad k_{j}\in[V_{s}]:\quad 1+\sum_{(k_{i},k_{j})\in E}1\geq\frac{|V_{s}|}{2\alpha}.$
By Algorithm 3 removes the vertex and its out-degree neighbors, which has the largest out-degree neighbors, in the remaining graph, the number of all vertices satisfies $|V_{s+1}|\leq(1-\frac{1}{2\alpha})|V_{s}|\leq|V_{s}|e^{-1/(2\alpha)}$.
When the graph $G$ has cycles, by Algorithm 3, given that $G_{s}=(V_{s},E)$ with the independence number $\alpha_{s}$, which is the final graph without cycles. We can discuss this in two cases.
@@
$|V_{s}|<\alpha$.
We can get an graph $G_{\eta}=(|V_{\eta}|,E)$ with $\eta\leq s$, which satisfies $|V_{\eta}|<\alpha$ and $|V_{\eta-1}|\geq\alpha$, since $V_{\eta+1}\leq Ke^{-1/(2\alpha)}$, we have $\eta-1\leq 2\alpha\ln(K/|V_{\eta}|)\leq 2\alpha\ln(K/\alpha)$. So we can guarantee
$|{\mathcal{D}}|\leq|V_{\eta}|+2\alpha\ln(K/\alpha)+1\leq\alpha(1+2\ln(K/\alpha)).$
@@
$|V_{s}|\geq\alpha$.
By $V_{s}\leq Ke^{-1/(2\alpha)}$, we have $s\leq 2\alpha\ln(K/|V_{s}|)\leq 2\alpha\ln(K/\alpha)$. Because $G_{s}$ has no cycles and the independence number $\alpha_{s}\leq\alpha$, we can get
$|{\mathcal{D}}|\leq\alpha_{s}+2\alpha\ln(K/\alpha)\leq\alpha(1+2\ln(K/\alpha)).$
The proof is complete.
B.2 Proof for Theorem 2
@@
[…]
@@
For each arm $k\not\in{\mathcal{A}}^{m}$, we have $n_{k}^{m}\leq\frac{16\lambda_{m}}{\Delta_{k}^{2}}$. Recalling the process of setting the actual expected number of pulls, for ease of analysis, if removing multiple arms from $H^{m}$ in one loop, we also consider this case as multi loops where the ${\mathcal{H}}_{s}^{m}$ is set by zero. Defining $b=K-|{\mathcal{A}}^{m}|$, by Lemma 12, we can set a parameter $\gamma^{m}=\max_{s=1,2,\cdots,b}|D_{s}^{m}|$. With loss of generality, for each arm $k\in{\mathcal{A}}^{m}$, we set $\Delta_{1}\geq\Delta_{2}\geq\cdots\geq\Delta_{b}$. Recalling the process of Algorithm 2, we only need to pull ${\mathcal{H}}_{s}^{m}$ times for all arms in the out-dominating set $D_{s}^{m}$, where we can observe at least ${\mathcal{H}}_{s}^{m}$ times for all arms $k\in[K]$. By this way, to maximize the value of $\sum_{k\in{\mathcal{A}}^{m}}R_{k}^{m}$, we should try to select the larger part of the suboptimal arm from the arms that have not been removed. In addition, we can have the following inequality:
${\mathcal{H}}_{1}^{m}\leq n_{1}^{m}=\frac{16\lambda_{m}}{\Delta_{1}^{2}},{\mathcal{H}}_{2}^{m}-{\mathcal{H}}_{1}^{m}\leq\frac{16\lambda_{m}}{\Delta_{2}^{2}},\cdots,{\mathcal{H}}_{b}^{m}-{\mathcal{H}}_{b-1}^{m}\leq\frac{16\lambda_{m}}{\Delta_{b}^{2}}.$
@@
$\begin{split}\sup\sum_{k\in{\mathcal{A}}^{m}}R_{k}^{m}&=\sup\sum_{s=1}^{b}{\mathcal{H}}_{s}^{m}\left(\sum_{k\in{\mathcal{A}}^{m},k\in D_{s}^{m}}\Delta_{k}\right)\\ &=\sum_{s=1}^{b-\gamma^{m}+1}\left(\frac{16\lambda_{m}}{\Delta_{s}^{2}}-\frac{16\lambda_{m}}{\Delta_{s-1}^{2}}\right)\left(\sum_{k=s}^{s+\gamma^{m}}\Delta_{k}\right)+\sum_{s=b-\gamma^{m}+2}^{b}\left(\frac{16\lambda_{m}}{\Delta_{s}^{2}}-\frac{16\lambda_{m}}{\Delta_{s-1}^{2}}\right)\left(\sum_{k=s}^{b}\Delta_{k}\right)\\ &=\sum_{k=1}^{\gamma^{m}}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{k=\gamma^{m}+1}^{b}\left(\frac{16\lambda_{m}}{\Delta_{k}^{2}}-\frac{16\lambda_{m}}{\Delta_{k-\gamma^{m}}^{2}}\right)\Delta_{k}\\ &\leq\sum_{k=\gamma^{m}+1}^{2\gamma^{m}}\frac{20\lambda_{m}}{\Delta_{k}}+\sum_{k=2\gamma^{m}+1}^{b}\left(\frac{16\lambda_{m}}{\Delta_{k}^{2}}-\frac{16\lambda_{m}}{\Delta_{k-1}^{2}}\right)\Delta_{k}\leq\cdots\leq\sum_{k=b-\gamma^{m}+1}^{b}\frac{32\lambda_{m}}{\Delta_{k}}.\end{split}$ (7)
@@
$=16\lambda_{m}\sum_{k=1}^{\gamma^{m}}\left(\frac{1}{\Delta_{k}}+\left(\frac{1}{\Delta_{\gamma^{m}+k}^{2}}-\frac{1}{\Delta_{k}^{2}}\right)\Delta_{\gamma^{m}+k}\right)\leq\sum_{k=\gamma^{m}+1}^{2\gamma^{m}}\frac{20\lambda_{m}}{\Delta_{k}}.$
@@
$\begin{split}{\mathbb{E}}[R(T)]&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(T)+3\log(T)\log(30K))}{\Delta_{k}}\\ &\quad+\frac{2^{12}K(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30K)))}{\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880C_{m}+1440+\sum_{\Delta_{k}>0}\frac{2^{14}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240K/\Delta))}{\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880C_{m}+1440+\sum_{k\in{\mathcal{I}}^{*}}\frac{2^{15}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240K/\Delta))}{\Delta}+\log(8/\Delta)\\ &=O\left(C+\sum_{k\in{\mathcal{I}}^{*}}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(K/\Delta)}{\Delta}\right).\end{split}$
Appendix C Proof of Theorem 3
C.1 Notations
Following [Zimmert et al., 2019, Dann et al., 2023], with loss of generality, we assume $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$ and $\Delta_{k}=\mu_{k}-\mu_{d}$ for all arms $k>d$.
C.2 Lemmas for Proving Theorem 3
@@
This proof is complete.
Using the methods mentioned before, all lemmas in Appendix A still holds. Maybe Lemma 7 will be different because we changed the way $r_{*}^{m}$ is assigned. Next we will prove that Lemma 7 still holds.
Proof 12
@@
$-\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}.$
@@
$r_{*}^{m}\leq\top(d)_{k}\left\{\mu_{k}+\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}\leq\mu_{d}+\frac{D_{m}}{N_{m}},$
$r_{*}^{m}=\top(d)_{k}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}\geq r_{d}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{d}^{m}}}\geq\mu_{d}-2\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{d}^{m}}}-\frac{D_{m}}{N_{m}},$
@@
$-\frac{D_{m}}{N_{m}}-2\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{d}^{m}}}\leq r_{*}^{m}-\mu_{d}\leq\frac{D_{m}}{N_{m}}.$
@@
Assuming the statement is true for the case of $m-1$, we have
@@
C.3 Proof for Theorem 3
Following to the work Zimmert et al. [2019], we can define the regret $R_{k}^{m}$ generated by arm $k$ in epoch $m$ as

### v2

@@ B.4.1 Lemmas for Proving Theorem 5
Lemma 20.
For any epoch $m$, all arms $k\neq k_{m}$ must satisfy $\tilde{n}_{k}^{m}\leq n_{k}^{m}$ and $\hat{n}_{k}^{m}\geq n_{k}^{m}$.
Proof.
@@
For each arm $k$ which does not have a self-loop, since $\widetilde{n}_{k_{m}}^{m}\geq 2^{-m}\geq n_{k}^{m}$ for all arms $k\neq k_{m}$, so we complete the proof. ∎
Same as the proof in Appendix B.1, we only need to change $\widetilde{n}_{k}^{m}$ to $\hat{n}_{k}^{m}$ to obtain the following lemmas.
Lemma 21.
@@
We can also always guarantee the following inequality:
@@
$\frac{3}{7}\Delta-\frac{3}{4}2^{-m}-6\rho_{m}\leq\Delta_{k^{*}}^{m}\leq\frac{1}{7}\Delta+2^{-(m-1)}+2\rho_{m}.$
Proof.
@@
$\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\hat{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\hat{n}_{k}^{m}}}.$
@@
$\frac{D_{m}}{N_{m}}-\frac{1}{4}\max\{\Delta_{k^{*}}^{m-1},\Delta^{m-1}\}\leq r_{*}^{m}-\mu_{k^{*}}\leq\frac{D_{m}}{N_{m}}.$
@@
Assuming the statement is true for m1m-1, for arm $k\neq k^{*}$, since there are at least two arms: one optimal arm and one sub-optimal arm, we can obtain
@@
This proof is complete. ∎
Lemma 24.
For any strongly observable directed graph $G$ with the independence number $\alpha$, the obtained out-domination set $D$ must satisfy $|D|\leq\lceil\alpha(1+2\ln(K/\alpha))\rceil$ by Algorithm 6. Especially, when $G$ is an acyclic graph and undirected graphs, we have $|D|\leq\alpha$.
Proof.
If the graph $G$ is undirected, then in each iteration of the Algorithm 6 we select a vertex and remove its neighbors. The selected vertices are therefore pairwise non-adjacent and form an independent set; hence the resulting out-dominating set $D$ satisfies $|D|\leq\alpha$. The remainder of our analysis focuses on the directed graphs.
Recalling the definition of the no-root vertex, for any strongly observable directed graph $G$ with the independence number $\alpha$, we can get that if we remove a no-root vertex $k$ and its out-degree neighbors, the remaining graph $G^{\prime}$ whose independence number is at most $\alpha-1$. That’s because no vertex in graph $G^{\prime}$ is connected to the vertex $k$. By the definition of strongly observable directed acyclic graph, we can always get the no-root vertex for the remaining graph, which means that $|D|\leq\alpha$.
Recalling the pseudo-code in Algorithm 6, we denote the loop index as $s$ and initialize $s=0$. In other words, $G_{s}=([V_{s}],E_{s})$ shows the residual graph obtained after $s$ loops of removing vertices. Given any graph $G_{s}=([V_{s}],E_{s})$ with no no-root vertices, $G_{s+1}$ is obtained by removing from $G_{s}$ the vertex $i_{s}$ with the largest out-degree $d_{s}^{+}$. Hence,
$d_{s}^{+}\geq\frac{|E_{S}|}{|V_{S}|}\geq\frac{|V_{s}|}{2\alpha_{s}}-\frac{1}{2}\geq\frac{|V_{s}|}{2\alpha}-\frac{1}{2},$
by Turan’s theorem (e.g., [alon2016probabilistic]), where $\alpha_{s}$ is the independence number of $G_{s}$ and $\alpha\geq\alpha_{s}$. This shows that
$|V_{s+1}|=|V_{s}|-d_{s}^{+}\leq|V_{s}|(1-\frac{1}{2\alpha})\leq|V_{s}|e^{-s/2\alpha}.$
Iterating, we obtain $|V_{s}|\leq Ke^{-1/2\alpha}$. We define the graph $G_{s_{0}}=([V_{s_{0}}],E_{s_{0}})$ with the independence number $\alpha_{s_{0}}$ as the the final residual graph has cycles. We can discuss this in two cases:
@@
$|V_{s_{0}}|<\alpha$.
We can get a graph $G_{s_{1}}=(|V_{s_{1}}|,E_{s_{1}})$ with $s_{1}\leq s_{0}$, which satisfies $|V_{s_{1}}|<\alpha$ and $|V_{s_{1}-1}|\geq\alpha$. Since $V_{s_{1}-1}\leq Ke^{-(s_{1}-1)/(2\alpha)}$, we have $s_{1}-1\leq\lceil 2\alpha\ln(K/|V_{s_{1}-1}|)\rceil\leq\lceil 2\alpha\ln(K/\alpha)\rceil$. So we can get
$|D|\leq|V_{s_{1}}|+\lceil 2\alpha\ln(K/\alpha)\rceil+1\leq\lceil\alpha(1+2\ln(K/\alpha))\rceil.$
@@
$|V_{s_{0}}|\geq\alpha$.
By $V_{s_{0}}\leq Ke^{-s_{0}/(2\alpha)}$, we have $s_{0}\leq\lceil 2\alpha\ln(K/|V_{s_{0}}|)\rceil\leq\lceil 2\alpha\ln(K/\alpha)\rceil$. Because $G_{s_{0}}$ has no cycles and the independence number $\alpha_{s_{0}}\leq\alpha$, we can get
$|D|\leq\alpha_{s_{0}}+\lceil 2\alpha\ln(K/\alpha)\rceil\leq\lceil\alpha(1+2\ln(K/\alpha))\rceil.$
The proof is complete. ∎
B.4.2 Proof for Theorem 5
@@
We define ${\mathcal{A}}^{m}\triangleq\left\{k\in[K]\,\big|\,0<\Delta_{k}\leq 64\rho_{m-1}\right\}$ for epoch $m$. By combining all three cases, we can upper bound the regret as
@@
Recalling the pseudo-code of SOG-BARBAT in Algorithm 4, we define the loop (Line 10-16) index as $a$. For ease of analysis, if removing multiple arms from $H^{m}$ in one loop, we also consider this case as multi loops where the $\overline{H}_{a}^{m}$ is set by zero. According to the above analysis, we have $n_{k}^{m}\leq\frac{16\lambda_{m}}{\Delta_{k}^{2}}$. Defining $b=K-|{\mathcal{A}}^{m}|$, by Lemma 24, we can set a parameter $\gamma^{m}=\max_{s=1,2,\cdots,b}|D_{a}^{m}|$. With loss of generality, for each arm $k\in{\mathcal{A}}^{m}$, we set $\Delta_{1}\geq\Delta_{2}\geq\cdots\geq\Delta_{b}$. Recalling the process of Algorithm 4, we only need to pull $\overline{H}_{a}^{m}$ times for all arms in the out-domination set $D_{s}^{m}$, where we can observe at least $\overline{H}_{a}^{m}$ times for all arms $k\in[K]$. By this way, to maximize the value of $\sum_{k\in{\mathcal{A}}^{m}}R_{k}^{m}$, we should try to select the larger part of the suboptimal arm from the arms that have not been removed. In addition, we can have the following inequality:
$\overline{H}_{1}^{m}\leq n_{1}^{m}=\frac{16\lambda_{m}}{\Delta_{1}^{2}},\overline{H}_{2}^{m}-\overline{H}_{1}^{m}\leq\frac{16\lambda_{m}}{\Delta_{2}^{2}},\cdots,\overline{H}_{b}^{m}-\overline{H}_{b-1}^{m}\leq\frac{16\lambda_{m}}{\Delta_{b}^{2}}.$
@@
$\begin{split}\sup\sum_{k\in{\mathcal{A}}^{m}}R_{k}^{m}&=\sup\sum_{s=1}^{b}\overline{H}_{a}^{m}\left(\sum_{k\in{\mathcal{A}}^{m},k\in D_{s}^{m}}\Delta_{k}\right)\\ &=\sum_{s=1}^{b-\gamma^{m}+1}\left(\frac{16\lambda_{m}}{\Delta_{s}^{2}}-\frac{16\lambda_{m}}{\Delta_{s-1}^{2}}\right)\left(\sum_{k=s}^{s+\gamma^{m}}\Delta_{k}\right)+\sum_{s=b-\gamma^{m}+2}^{b}\left(\frac{16\lambda_{m}}{\Delta_{s}^{2}}-\frac{16\lambda_{m}}{\Delta_{s-1}^{2}}\right)\left(\sum_{k=s}^{b}\Delta_{k}\right)\\ &=\sum_{k=1}^{\gamma^{m}}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{k=\gamma^{m}+1}^{b}\left(\frac{16\lambda_{m}}{\Delta_{k}^{2}}-\frac{16\lambda_{m}}{\Delta_{k-\gamma^{m}}^{2}}\right)\Delta_{k}\\ &\leq\sum_{k=\gamma^{m}+1}^{2\gamma^{m}}\frac{20\lambda_{m}}{\Delta_{k}}+\sum_{k=2\gamma^{m}+1}^{b}\left(\frac{16\lambda_{m}}{\Delta_{k}^{2}}-\frac{16\lambda_{m}}{\Delta_{k-1}^{2}}\right)\Delta_{k}\leq\cdots\leq\sum_{k=b-\gamma^{m}+1}^{b}\frac{32\lambda_{m}}{\Delta_{k}}.\end{split}$ (16)
@@
$\qquad\qquad\qquad\qquad=16\lambda_{m}\sum_{k=1}^{\gamma^{m}}\left(\frac{1}{\Delta_{k}}+\left(\frac{1}{\Delta_{\gamma^{m}+k}^{2}}-\frac{1}{\Delta_{k}^{2}}\right)\Delta_{\gamma^{m}+k}\right)\leq\sum_{k=\gamma^{m}+1}^{2\gamma^{m}}\frac{20\lambda_{m}}{\Delta_{k}}.$
@@
$\begin{split}R(T)&\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{m=1}^{M}\sum_{\Delta_{k}>0}\frac{16\lambda_{m}}{\Delta_{k}}+\sum_{m=1}^{\log_{2}(8/\Delta)}\left(\frac{16K\lambda_{m}}{\Delta}+1\right)\\ &\leq\sum_{m=1}^{M-1}1440{\mathbb{E}}[D_{m}]+\sum_{\Delta_{k}>0}\frac{2^{12}(\log^{2}(T)+3\log(T)\log(30K))}{\Delta_{k}}\\ &\quad+\frac{2^{12}K(\log^{2}(8/\Delta)+3\log(8/\Delta)\log(30K)))}{\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880C_{m}+1440+\sum_{\Delta_{k}>0}\frac{2^{14}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240K/\Delta))}{\Delta}+\log(8/\Delta)\\ &\leq\sum_{m=1}^{M-1}2880C_{m}+1440+\sum_{k\in{\mathcal{I}}^{*}}\frac{2^{15}\log(T)\log(30KT)}{\Delta_{k}}\\ &\quad+\frac{2^{14}K\log(8/\Delta)\log(240K/\Delta))}{\Delta}+\log(8/\Delta)\\ &=O\left(C+\sum_{k\in{\mathcal{I}}^{*}}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(K/\Delta)}{\Delta}\right).\end{split}$
where $\mathcal{I}$$^{*}$ is the set of at most $\lfloor\alpha(1+2\ln(K/\alpha))\rfloor$ arms with the smallest gaps. Especially, for directed acyclic graphs (including undirected graphs), $\mathcal{I}$$^{*}$ is the set of at most $\alpha$ arms with the smallest gaps.
B.5 Proof of Theorem 6
B.5.1 Notations
Following [zimmert2019beating, dann2023blackbox], with loss of generality, we assume $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$ and $\Delta_{k}=\mu_{k}-\mu_{d}$ for all arms $k>d$.
B.5.2 Lemmas for Proving Theorem 6
@@
This proof is complete. ∎
Using the methods mentioned before, all lemmas in Appendix B.1 still holds. Maybe Lemma 7 will be different because we changed the way $r_{*}^{m}$ is assigned. Next we will prove that Lemma 7 still holds.
Proof.
@@
$\frac{D_{m}}{N_{m}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\leq r_{k}^{m}-\mu_{k}\leq\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}.$
@@
$r_{*}^{m}\leq\top_{d}\left(\left\{\mu_{k}+\frac{D_{m}}{N_{m}}+\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}_{k\in[K]}\right)\leq\mu_{d}+\frac{D_{m}}{N_{m}},$
$r_{*}^{m}=\top_{d}\left(\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}_{k\in[K]}\right)\geq r_{d}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{d}^{m}}}\geq\mu_{d}-2\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{d}^{m}}}-\frac{D_{m}}{N_{m}},$
@@
$\frac{D_{m}}{N_{m}}-2\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{d}^{m}}}\leq r_{*}^{m}-\mu_{d}\leq\frac{D_{m}}{N_{m}}.$
@@
Assuming the statement is true for the case of m1m-1, we have
@@
where the third inequality comes from the upper bound of $\Delta_{k}^{m-1}$. ∎
B.5.3 Proof for Theorem 6
Following to the work [zimmert2019beating], we can define the regret $R_{k}^{m}$ generated by arm $k$ in epoch $m$ as